Gianna:

Milo bought 2 1/2 pounds of red apples and 3 3/4 pounds of green apples to make applesauce. How many pounds of apples did he buy in all? a. Write an expression that models the problem. b. What is the LCD of the fractions in your expression? Explain how you found the LCD. c. Evaluate the expression. d. Answer the question asked in the problem.

jhonyy9:

let be x the red apples and y the green apples so in this case you can writing x = 2 1/2 pounds and y = 3 3/4 pounds x = 2 1/2 = (2*2 +1)/2 = 5/2 and y = 3 3/4 = (3*4+3)/4 = 15/4 a). x +y = ? b). bc. denominators are 2 and 4 so how you get the LCD - (low common denominator ) ? hope helped